OpenStudy (anonymous):

Find the solution of this system of equations. separate the x-and y- values with a comma. x-2y=-26 x-y=-2

imqwerty (imqwerty):

x-2y=-26 x-y=-2 subttract the 2nd equation frm the 1st one then u'll get the value of y. then plug it into any of the equations to get value of x

OpenStudy (anonymous):

I tried that already and it was a wrong answer

OpenStudy (anonymous):

Steps: I will try to solve your system of equations. x−2y=−26;x−y=−2 Step: Solve x−2y=−26 for x: x−2y+2y=−26+2y(Add 2y to both sides) x=2y−26 Step: Substitute (2y−26) for x in x−y=−2: x−y=−2 2y−26−y=−2 y−26=−2(Simplify both sides of the equation) y−26+26=−2+26(Add 26 to both sides) y=24 Step: Substitute (24) for y in x=2y−26: x=2y−26 x=(2)(24)−26 x=22(Simplify both sides of the equation) Answer: x=22 and y=24

OpenStudy (anonymous):

x+6y= 23x-y=5 is the one i am doing now because the last one was wrong ....

OpenStudy (anonymous):

and thank you

OpenStudy (anonymous):

yes